### v1

# AsEP: Benchmarking Deep Learning Methods for Antibody-specific Epitope Prediction

Abstract
Epitope identification is vital for antibody design yet challenging due to the inherent variability in antibodies. While many deep learning methods have been developed for general protein binding site prediction tasks, whether they work for epitope prediction remains an understudied research question. The challenge is also heightened by the lack of a consistent evaluation pipeline with sufficient dataset size and epitope diversity. We introduce a filtered antibody-antigen complex structure dataset, AsEP (Antibody-specific Epitope Prediction). AsEP is the largest of its kind and provides clustered epitope groups, allowing the community to develop and test novel epitope prediction methods. AsEP comes with an easy-to-use interface in Python and pre-built graph representations of each antibody-antigen complex while also supporting customizable embedding methods. Based on this new dataset, we benchmarked various representative general protein-binding site prediction methods and find that their performances are not satisfactory as expected for epitope prediction. We thus propose a new method, WALLE, that leverages both protein language models and graph neural networks. WALLE demonstrate about $5$X performance gain over existing methods. Our empirical findings evidence that epitope prediction benefits from combining sequential embeddings provided by language models and geometrical information from graph representations, providing a guideline for future method design. In addition, we reformulate the task as bipartite link prediction, allowing easy model performance attribution and interpretability. We open-source our data and code at https://github.com/biochunan/AsEP-dataset.
1 Introduction
Antibodies are specialized proteins produced by our immune system to combat foreign substances called antigens. Their unique ability to bind with high affinity and specificity sets them apart from regular proteins and small-molecule drugs, making them increasingly popular in therapeutic engineering.
While the community is shifting toward computational antibody design given a pre-determined epitope  (Jin et al., 2022; Zhou et al., 2024; Bennett et al., 2024), accurate prediction of epitopes remains underexplored. Accurate identification of epitopes is beneficial for understanding antibody-antigen interactions, antibody function, and streamlining antibody engineering. This task remains challenging due to multiple factors (Akbar et al., 2022; Hummer et al., 2022), such as, the lack of comprehensive datasets, limited interpretability and generalizability. Available datasets are limited in size, up to 582 complexes from Bepipred-3.0 (Clifford et al., 2022), and feature disproportionate representation among various epitopes.
Existing methods perform poorly on epitope prediction task (Cia et al., 2023), with a ceiling MCC (Matthew’s Correlation Coefficient) of $0.06$. However, recent advancements in graph-based learning, coupled with an increase in available antibody structures in the Protein Data Bank (PDB) (Berman et al., 2003), highlight the need to reevaluate current methods and establish a benchmark dataset for predicting antibody-antigen interactions.
We approach the problem as a bipartite graph link prediction task. Traditionally, graph link prediction focuses on identifying connections within the same graph, such as in protein-protein interaction networks. Our research extends this concept to bipartite graphs at the molecular level and proposes our own model, WALLE. Because existing methods generally predict protein binding sites or epitopes in antibody-antigen complexes rather than residue-residue interactions, we focus on benchmarking the node classification task while providing WALLE’s performance on the bipartite link prediction task as a baseline.
2 Previous Work
Accurate epitope prediction for antibody design remains challenging due to the complexity of antibody-antigen interactions and limitations of existing datasets and methods. Several computational approaches have been developed, but they often fall short in terms of accuracy and applicability. Here, we present a representative set of state-of-the-art methods.
EpiPred (Krawczyk et al., 2014) implements a graph-based antibody-antigen specific scoring function that considers all possible residue-residue pairs at the interface between anybody and antigen structures. It samples surface patches from the antigen structure surface and selects the highest-ranked patch with the highest score as the predicted set of epitope residues.
ESMFold(Lin et al., 2023) is a protein language model based on ESM2 (Lin et al., 2023), and its folding head was trained on over 325 thousand protein structures. It achieves comparable performance to AlphaFold2 (Jumper et al., 2021) and is included in our benchmarking due to its faster processing.
We also include methods that consider only antigens:
ESMBind (Schreiber, 2023) is a language model that predicts protein binding sites for single protein sequence inputs. It is fine-tuned based on ESM2 (Lin et al., 2023) using Low-Rank Adaptation (Hu et al., 2021) on a dataset composed of more than 200 thousand protein sequences with annotated binding sites.
MaSIF-site (Gainza et al., 2020) is a geometric deep learning method that predicts binding sites on the surface of an input protein structure. It converts antigen surfaces into mesh graphs, with each mesh vertex encoded with geometric and physicochemical descriptors, and predicts binding sites as a set of mesh vertices.
PECAN and EPMP (Pittala & Bailey-Kellogg, 2020; Vecchio et al., 2021) are graph neural networks that predict epitope residues taken antibody-antigen structure pairs as input. They use position-specific scoring matrices (PSSM) as node embeddings and graph attention networks to predict the binary labels of nodes in the antigen graph. For details, we refer readers to  Section A.1.
Two Complementary Surveys are notable:
Zhao et al. (2024) benchmarked docking methods like ZDOCK (Pierce et al., 2011), ClusPro (Kozakov et al., 2017), and HDOCK (Yan et al., 2017), and Alphafold-Multimer (Evans et al., 2022) on a set of $112$ antibody-antigen complexes. They showed that all docking methods gave a success rate of $8.0\%$ at most if using the top 5 decoys; Alphafold-Multimer showed a better performance with a $15.3\%$ success rate.
Cia et al. (2023) focused on epitope prediction using a dataset of $268$ complexes, defining epitope residues as having at least a $5\%$ change in relative solvent accessibility upon complex formation. They benchmarked various methods, finding existing methods insufficient for accurate epitope prediction.
3 Problem Formulation
Antibody-antigen interaction is important for analyzing protein structures. The problem can be formulated as a bipartite graph link prediction task. The inputs are two disjoint graphs, an antibody graph $G_{A}=\left(V_{A},E_{A}\right)$ and an antigen graph $G_{B}=\left(V_{B},E_{B}\right)$, where $V_{x}$ is the vertice set for graph $x$ and $E_{x}$ is the edge set for graph $x$. Since the neural networks only take continuous values as input, we also have a function $h$ to encode each vertex into a vector $h:V\to\mathbb{R}^{D}$. The design of the encoding function depends on the methods. For example, $h$ can be a one-hot encoding layer or pretrained embeddings given by a protein language model. We use different encoding functions for antibodies and antigens: $h_{A}:V_{A}\to\mathbb{R}^{D_{A}}$, and $h_{B}:V_{B}\to\mathbb{R}^{D_{B}}$.
In addition, $E_{A}\in\{0,1\}^{|V_{A}|\times|V_{A}|}$ and $E_{B}\in\{0,1\}^{|V_{B}|\times|V_{B}|}$ denote the adjacency matrices for the antibody and antigen graphs, respectively. In this work, the adjacency matrices are calculated based on the distance matrix of the residues. Each entry $e_{ij}$ denotes the proximity between residue $i$ and residue $j$; $e_{ij}=1$ if the Euclidean distance between any non-hydrogen atoms of residue $i$ and residue $j$ is less than $4.5$Å, and $e_{ij}=0$ otherwise. The antibody graph $G_{A}$ is constructed by combining the CDR residues from the heavy and light chains of the antibody, and the antigen graph $G_{B}$ is constructed by combining the surface residues of the antigen. The antibody and antigen graphs are disjoint, i.e., $V_{A}\cap V_{B}=\emptyset$.
Figure 1: An example to illustrate interacting residues. The two dashed lines denote the distance between non-hydrogen atoms from different interacting residues from two different protein chains.
We consider two subtasks based on these inputs.
Epitope Prediction Epitopes are the regions on the antigen surface recognized by antibodies; in other words, they are a set of antigen residues in contact with the antibody and are determined from the complex structures using the same distance cutoff of $4.5$Å as aforementioned. For a node in the antigen graph $v\in V_{B}$, if there exists a node in the antibody graph $u\in V_{A}$ such that the distance between them is less than $4.5$Å, then $v$ is an epitope node. Epitope nodes and the remaining nodes in $G_{B}$ are assigned labels of $1$ and 00, respectively. The first task is then a node classification within the antigen graph $G_{B}$ given the antibody graph $G_{A}$.
This classification takes into account the structure of the antibody graph, $G_{A}$, mirroring the specificity of antibody-antigen binding interactions. Different antibodies can bind to various antigen locations, corresponding to varying subsets of epitope nodes in $G_{B}$. This question differs from conventional epitope prediction methods that do not consider the antibody structure and end up predicting the likelihood of the subset of antigen nodes serving as epitopes, such as ScanNet (Tubiana et al., 2022), MaSIF (Gainza et al., 2020).
The task is to develop a binary classifier $f:V_{B}\to\{0,1\}$ that takes both the antibody and antigen graphs as input and predicts the label for antigen nodes as formulated below:
$f(v\,;G_{B},G_{A})$ =$=\begin{cases}1&\text{if }v\text{ is an epitope};\\ 0&\text{otherwise}.\end{cases}$ (1)
Bipartite Link Prediction The second task takes it further by predicting concrete interactions between nodes in $G_{A}$ and $G_{B}$, resulting in a bipartite graph that represents these antibody-antigen interactions. Moreover, this helps attribute the model performance to specific interactions at the molecular level and provide more interpretability. Accurately predicting these interactions is critical for understanding the binding mechanisms and for guiding antibody engineering.
We model the antibody-antigen interaction as a bipartite graph:
$K_{m,n}=(V_{A},V_{B},E)$
where $m=|V_{A}|$ and $n=|V_{B}|$ denote the number of nodes in the two graphs, respectively, and $E$ denotes all possible inter-graph links. In this bipartite graph, a node from the antibody graph is connected to each node in the antigen graph via an edge $e\in E$. The task is then to predict the label of each bipartite edge. If the residues of a pair of nodes are located within $4.5$Å of each other, referred to as in contact, the edge is labeled as $1$; otherwise, 00. For any pair of nodes, denoted as $(v_{a},v_{b})\ \forall v_{a}\in V_{A},v_{b}\in V_{B}$, the binary classifier $g:K_{m,n}\rightarrow\{0,1\}$ is formulated as below:
$g(v_{a},v_{b};K_{m,n})=\begin{cases}1&\text{if $v_{a}$ and $v_{b}$ are in contact}\\ 0&\text{otherwise}.\end{cases}$ (2)
4 AsEP dataset
We present our dataset AsEP of filtered, cleaned and processed antibody-antigen complex structures. It is the largest collection of antibody-antigen complex structures to our knowledge. Antibodies are composed of two heavy chains and two light chains, each of which contains a variable domain (areas of high sequence variability) composed of a variable heavy (VH) and a variable light (VL) domain responsible for antigen recognition and binding (Chothia & Lesk, 1987). These domains have complementarity-determining regions (CDR, Figure 2 top blue, yellow, and red regions), which are the primary parts of antibodies responsible for antigen recognition and binding.
Figure 2: Graph visualization of an antibody-antigen complex. Top: the molecular structure of an antibody complexed with the receptor binding domain of SARS-Cov-2 virus (PDB code: 7KFW), the antigen. Spheres indicate the alpha carbon atoms of each amino acid. Color scheme: the antigen is colored in magenta, the framework region of the heavy and light chains is colored in green and cyan and CDR 1-3 loops are colored in blue, yellow, and red, respectively. Bottom: the corresponding graph. Green vertices are antibody CDR residues and pink vertices are antigen surface residues.
4.1 Antibody-antigen complexes
We sourced our initial dataset from the Antibody Database (AbDb) (Ferdous & Martin, 2018), dated 2022/09/26, which contains $11,767$ antibody files originally collected from the Protein Data Bank (PDB) (Berman et al., 2003). We extracted conventional antibody-antigen complexes that have a VH and a VL domain with a single-chain protein antigen, and there are no unresolved CDR residues due to experimental errors, yielding $4,081$ antibody-antigen complexes. To ensure data balance, we removed identical complexes using an adapted version of the method described in Krawczyk et al. (2014). We clustered the complexes by antibody heavy and light chains followed by antigen sequences using MMseqs2 (Steinegger & Söding, 2017). We retained only one representative complex for each unique cluster, leading to a refined dataset of $1,725$ unique complexes. Two additional complexes were manually removed; CDR residues in the complex 6jmr_1P are unknown (labeled as ‘UNK’) and it is thus impossible to build graph representations upon this complex; 7sgm_0P was also removed because of non-canonical residues in its CDR loops. The final dataset consists of $1,723$ antibody-antigen complexes. For detailed setup and processing steps, please refer to Section A.3.
4.2 Convert antibody-antigen complexes to graphs
These $1,723$ files were then converted into graph representations, which are used as input for WALLE. In these graphs, each protein residue is modeled as a vertex. Edges are drawn between pairs of residues if any of their non-hydrogen atoms are within $4.5$Å of each other, adhering to the same distance criterion used in PECAN (Pittala & Bailey-Kellogg, 2020).
Exclude buried residues In order to utilize structural information effectively, we focused on surface residues, as only these can interact with another protein. Consequently, we excluded buried residues, those with a solvent-accessible surface area of zero, from the antigen graphs. The solvent-accessible surface areas were calculated using DSSP (Kabsch & Sander, 1983) via Graphein (Jamasb et al., 2021). It is important to note that the number of interface nodes are much smaller than the number of non-interface nodes in the antigen, making the classification task more challenging.
Exclude non-CDR residues We also excluded non-CDR residues from the antibody graph, as these are typically not involved in antigen recognition and binding. This is in line with the approach adopted by PECAN (Pittala & Bailey-Kellogg, 2020) and EPMP (Vecchio et al., 2021)Figure 2 provides a visualization of the processed graphs.
Node embeddings To leverage the state-of-the-art protein language models, we generated node embeddings for each residue in the antibody and antigen graphs using AntiBERTy (Ruffolo et al., 2021) (via IgFold (Ruffolo et al., 2023) package) and ESM2 (Lin et al., 2022) (esm2_t12_35M_UR50D) models, respectively. In our dataset interface package, we also provide a simple embedding method using one-hot encoding for amino acid residues. Other node embedding methods can be easily incorporated into our dataset interface.
4.3 Dataset split
We propose two types of dataset split settings. The first is a random split based on the ratio of epitope to antigen surface residues, #epitope nodes#antigen $\frac{\#\text{epitope nodes}}{\#\text{antigen nodes}}$; the second is a more challenging setting where we split the dataset by epitope groups. The first setting is straightforward and used by previous methods. The second setting is more challenging because it requires the model to generalize to unseen epitope groups.
Split by epitope to antigen surface ratio As aforementioned, the number of non-interface nodes in the antigen graph is much larger than the number of interface nodes. While epitopes usually have a limited number of residues, typically around $14.6\pm 4.9$ amino acids (Reis et al., 2022), the antigen surface may extend to several hundred or more residues. The complexity of the classification task therefore increases with the antigen surface size. To ensure similar complexity among train, validation, and test sets, we stratified the dataset to include a similar distribution of epitope to non-epitope nodes in each set. Table S3 shows the distribution of epitope-to-antigen surface ratios in each set. This led to $1383$ antibody-antigen complexes for the training set and $170$ complexes each for the validation and test sets. The list of complexes in each set is provided in the Supplementary Table SI-split-epitope-ratio.csv.
Split by epitope groups This is motivated by the fact that antibodies are highly diverse in the CDR loops and by changing the CDR sequences it is possible to engineer novel antibodies to bind different sites on the same antigen. This was previously observed in the EpiPred dataset where Krawczyk et al. (2014) tested the specificity of their method on five antibodies associated with three epitopes on the same antigen, hen egg white lysozyme.
We inlcude $641$ unique antigens and $973$ epitope groups in our dataset. We include multi-epitope antigens. For example, there are $64$ distinct antibodies that bind to coronavirus spike protein.
We can see that different antibodies bind to different locations on the same antigen. Details of all epitope groups are provided in the Supplementary Table SI-AsEP-entries.csv. We then split the dataset into train, validation, and test sets such that the epitopes in the test set are not found in either train or validation sets. We followed an 80%/10%/10% split for the number of complexes in each set. This resulted in $1383$ complexes for the training set and $170$ complexes for the validation and test sets. The list of complexes in each set is provided in the Supplementary Table SI-split-epitope-group.csv.
User-friendly Dataset Interface We implemented a Python package interface for our dataset using PyTorch Geometric (Fey & Lenssen, 2019). Users can load the dataset as a PyTorch Geometric dataset object and use it with PyTorch Geometric’s data loaders. We provide an option to load node embeddings derived from AntiBERTy and ESM2 or simply one-hot embeddings. Each data object in the dataset is a pair of antibody and antigen graphs; both node- and edge-level labels are provided, and the node-level labels are used for the epitope prediction task.
5 WALLE: a graph-based method for epitope prediction
Alongside our dataset interface, we also provide a graph-based model named WALLE. It takes as input a pair of antibody and antigen graphs, constructed as detailed above for the AsEP dataset, and makes node- and edge-level predictions.
Figure 3: A schematic of the preprocessing step that turns an input antibody-antigen complex structure into a graph pair and the model architecture of WALLE.
Graph Modules The architecture of WALLE incorporates graph modules that process the input graphs of antibody and antigen structures, as depicted in Figure 3). Inspired by PECAN and EPMP, our model treats the antibody and antigen graphs separately, with distinct pathways for each. The antibody graph is represented by node embeddings $X_{A}$ with a shape of $(M,512)$ and an adjacency matrix $E_{A}$, while the antigen graph is described by node embeddings $X_{B}$ with a shape of $(N,480)$ and its corresponding adjacency matrix $E_{B}$. The embedding sizes are consistent with AntiBERTy and ESM2 (esm2_t12_35M_UR50D).
Both antibody and antigen graph nodes are first projected into the dimensionality of $128$ using fully connected layers. The resulting embeddings are then passed through two Graph Convolutional Network (GCN) modules consecutively to refine the features and yield updated node embeddings $X_{A}^{\prime}$ and $X_{B}^{\prime}$ with a reduced dimensionality of $(M,64)$. The output from the first GCN layer is passed through a ReLU activate function. Outputs from the second GCN layer are directly fed into the Decoder module. These GCNs operate independently, each with its own parameters, ensuring that the learned representations are specific to the antibody or the antigen.
The use of separate GCN modules for the antibody and antigen allows for the capture of unique structural and functional characteristics pertinent to each molecule before any interaction analysis. This design choice aligns with the understanding that the antibody and antigen have distinct roles in their interactions and their molecular features should be processed separately.
Decoder
We used a simple decoder to predict the binary labels of edges between the antibody and antigen graphs, which takes a pair of node embeddings output by the graph modules as input and predicts the probability of each edge. An edge is assigned a binary label of $1$ if the predicted probability is greater than $0.5$ or 00 otherwise. This is shown as the Decoder module in Figure 3. For the epitope prediction task, we convert edge-level predictions to node-level by summing the predicted probabilities of all edges connected to an antigen node; we assign the antigen node a label of $1$ if the number of connected edges is greater than a threshold or 00 otherwise. The threshold is treated as a hyperparameter and is optimized in the experiments.
Implementation
We used PyTorch Geometric (Fey & Lenssen, 2019) framework to build our model. The graph modules are implemented using the GCNConv module from PyTorch Geometric. We trained the model to minimize a loss function consisting of two parts: a weighted binary cross-entropy loss for the bipartite graph link reconstruction and a regularizer for the number of positive edges in the reconstructed bipartite graph. We used the same set of hyperparameters and loss functions for both dataset settings. The loss function and hyperparameters are described in detail in Section A.6.
Table 1: Performance on test set from dataset split by epitope to antigen surface ratio and epitope groups.
(a) Performance on dataset split by epitope to antigen surface ratio.
Algorithm MCC Precision Recall AUCROC F1
WALLE 0.210 (0.020) 0.235 (0.018) 0.422 (0.028) 0.635 (0.013) 0.258 (0.018)
EpiPred 0.029 (0.018) 0.122 (0.014) 0.180 (0.019) 0.142 (0.016)
ESMFold 0.028 (0.010) 0.137 (0.019) 0.043 (0.006) 0.060 (0.008)
ESMBind 0.016 (0.008) 0.106 (0.012) 0.121 (0.014) 0.506 (0.004) 0.090 (0.009)
MaSIF-site 0.037 (0.012) 0.125 (0.015) 0.183 (0.017) 0.114 (0.011)
(b) Performance on dataset split by epitope groups.
Algorithm MCC Precision Recall AUCROC F1
WALLE 0.077 (0.015) 0.143 (0.017) 0.266 (0.025) 0.544 (0.010) 0.145 (0.014)
EpiPred -0.006 (0.015) 0.089 (0.011) 0.158 (0.019) 0.112 (0.014)
ESMFold 0.018 (0.010) 0.113 (0.019) 0.034 (0.007) 0.046 (0.009)
ESMBind 0.002 (0.008) 0.082 (0.011) 0.076 (0.011) 0.500 (0.004) 0.064 (0.008)
MaSIF-site 0.046 (0.014) 0.164 (0.020) 0.174 (0.015) 0.128 (0.012)
MCC: Matthews Correlation Coefficient; AUCROC: Area Under the Receiver Operating Characteristic Curve; F1: F1 score. Standard errors are included in the parentheses. We omitted the results of EpiPred, ESMFold and MaSIF-site for AUCROC. For EpiPred and ESMFold, the interface residues are determined from the predicted structures by these methods such that the predicted values are binary and not comparable to other methods; As for MaSIF-site, it outputs the probability of mesh vertices instead of node probabilities and epitopes are determined as residues close to mesh vertices with probability greater than 0.7.
Table 2: Summary of Features Used in Benchmarking Methods.
Antibody Structure PLM Graph
WALLE
EpiPred $\times$
ESMFold $\times$ $\times$
MaSIF-site $\times$ $\times$
ESMBind $\times$ $\times$ $\times$
Antibody: Antibody is taken into consideration when predicting epitope nodes;
Structure: Topological information from protein structures;
PLM: Representation from Protein Language Models;
Graph: Graph representation of protein structures.
Experiment results We evaluated each method for both dataset split settings on the test set using the metrics described in Section A.2. Table 1(a) and Table 1(b) summarize the average performance metrics across the test set samples. WALLE generally shows better performance among all metrics except for recall for both dataset splits. We also provided the baseline performance for bipartite link prediction in Table S6.
We also carried out ablation studies (Appendix C) to investigate the impact of different components of WALLE. When we replace the graph convolutional layers with fully connected layers, its performance degenerates considerably, suggesting that the graph convolutional layers contribute to the model’s performance. This is related to the fact that the interaction between a pair of protein structures is dependent on the spatial arrangement of the residues as discussed by Reis et al. (2022). The interface polar bonds, a major source of antibody specificity, tend to shield interface hydrophobic clusters. In addition, the language model embeddings also contribute to the model’s performance, as performance drops when they are replaced by one-hot or BLOSUM62 embeddings. Finally, we also investigated whether the choice of language model affects the model’s performance. We found that the model using AntiBERTy and ESM2 embeddings for antibodies and antigens performed slightly better than the model using ESM2 embeddings for both antibodies and antigens. This suggests that the choice of language model may impact the model’s performance, but a model like ESM2, which is trained on general protein sequences, may contain sufficient information for the epitope prediction task.
While WALLE outperforms other methods in the second dataset split setting, its performance degenerated considerably from the first dataset split setting. This suggests that WALLE is likely biased toward the epitopes in the training set and does not generalize well to unseen epitopes. The performance of the other four methods is not ideal for this task. We believe this would require a more sophisticated model architecture and a more comprehensive dataset to improve the performance of epitope prediction. We invite researchers to examine this issue and bring their insights to improve generalization.
Edge features In terms of structure representation, we only used a simple invariant edge feature, the distance matrix, to capture the neighborhood information of each residue. This topological descriptor already performs better than other methods that use sequence-based features. For future work, more edge features can be incorporated to enrich the graph representation, in addition to the invariant edge features used in this work, such as inter-residue distances and edge types used in GearNet (Zhang et al., 2023), and SE3 equivariant features, such as rotational and orientational relationships between residues as used in abdockgen (Jin et al., 2022).
Antibody types We also plan to extend our work to include other types of antibodies. Current work only looks at conventional antibodies, consisting of heavy- and light-chain variable domains. There are also an increasing number of novel antibodies, for example nanobodies, which are single-variable-domain antibodies derived from camelids. These will be included in future work.
6 Conclusion
In this work, we proposed a novel benchmarking dataset for the epitope prediction task and clustered the samples by epitopes. We also provided a model, WALLE, which combines protein language models and graph neural networks to leverage their abilities in capturing amino acid contextual and geometric information. We benchmarked WALLE and four other methods, showing that while WALLE outperforms existing methods on both tasks, there remains room for improvement. Nevertheless, we showed the combination of protein language models and graph neural networks improved the performance on the epitope prediction task. We also discussed possible future directions. This work can serve as a starting point for future research in this area.
7 Acknowledgments
CL was part-funded by a UCL Centre for Digital Innovation Amazon Web Services (AWS) Scholarship; LD was funded by an ISMB MRC iCASE Studentship (MR/R015759/1).
Appendix A Appendix-A
A.1 Related work
Comparison of Previous Datasets We would like to highlight our dataset, AsEP, is the largest curated AbAg benchmarking dataset to date. Existing ones either focus on general protein-protein complexes designed to develop general docking methods or are way smaller than AsEP if designed for AbAg interaction research. We summarized the sizes of existing datasets in the following table.
Table S1: Comparison of Dataset Sizes Across Different Methods
Method Dataset Size
WALLE (AsEP) 1723 AbAg complexes
Gao et al. 2022 Gao et al. (2022) 258 AbAg complexes
CSM-AB Myung et al. (2021) 472 AbAg complexes
SAGERank Sun et al. (2023) 287 AbAg complexes
Bepipred3.0 Clifford et al. (2022) 582 AbAg complexes
SCEptRe Mahajan et al. (2019) is a related dataset that keeps a weekly updated collection of 3D complexes of epitope and receptor pairs, for example, antibody-antigen, TCR-pMHC, and MHC-ligand complexes derived from the Immune Epitope Database (IEDB). Our approach for clustering antibody-antigen complexes regarding their epitopes is similar to theirs, with the difference in the clustering strategy. We cluster by antigen, then epitope group, and we allow mutated amino acids in the same epitope region because we turn the epitope sites into columns in the multiple sequence alignment. In contrast, SCEptRe clusters by antibody and then compares epitope similarity by epitope conformation using atom-pair distances via PocketMatch Yeturu & Chandra (2008), which is beneficial for comparing the function of various paratopes but is less suitable for our task of predicting epitope residues.
Sequence-based epitope predictor We also tested purely sequence-based epitope prediction tool, for example, Bepipred3.0 Clifford et al. (2022) on our dataset. Bepipred3.0 uses ESM2 model, esm2_t33_650M_UR50D to generate sequence embeddings and was trained on a smaller dataset of 582 antibody-antigen structures and evaluated on 15 antibody-antigen complexes. The authors provided a relatively larger evaluation of linear B-cell epitopes derived from the Immune Epitope Database and reported an AUC-ROC of $0.693$ on the top $10\%$ of the predicted epitopes. We tested Bepipred3.0 on our dataset and found its performance degenerates significantly, as shown in the table below. This is not surprising because linear epitopes are consecutive positions in an antigen sequence, and this task fits better with language model design. Additionally, as pointed out by the authors, approximately $90\%$ of epitopes (B-cell) fall into the conformational category Clifford et al. (2022), which highlights the importance of the present benchmark dataset composed of conformational epitopes derived from filtered antibody-antigen  structures. We believe these results underline the findings in our paper, showing that large language models alone, even if specialized for antibody-antigen interactions, do not encompass all the relevant information needed for epitope prediction.
Confidence Threshold Top 10% Top 30% Top 50% Top 70% Top 90%
AUC 0.693392 0.693392 0.693392 0.693392 0.693392
Balanced Accuracy 0.573132 0.636365 0.638755 0.604556 0.542274
MCC 0.109817 0.140183 0.134689 0.113372 0.071876
Precision-Recall AUC 0.176429 0.176429 0.176429 0.176429 0.176429
Accuracy 0.850178 0.701202 0.536947 0.362489 0.179051
Precision 0.169202 0.141361 0.120547 0.104286 0.090441
Recall 0.236760 0.553204 0.756607 0.892628 0.977723
F1-Score 0.173370 0.208366 0.197153 0.179294 0.160151
Table S2: Bepipred3.0 results for the presented AsEP dataset. The distance cutoff was changed to 4.0Å, as this is the threshold used by Bepipred3.0. Results are shown for five confidence thresholds as described in the BepiPred-3.0 paper. Across all stringency settings and metrics, Bepipred scored lower than Walle. Furthermore, it is possible that some of the structures within the dataset are contained within the Bepipred3.0 dataset, artificially increasing scores.
A.2 Evaluation
In this work, we focus on the epitope prediction task. We evaluate the performance of each method using consistent metrics. Matthew’s Correlation Coefficient (MCC) is highly recommended for binary classification assessments Matthews (1975) and is especially advocated for its ability to provide equal weighting to all four values in the confusion matrix, making it a more informative metric about the classifier’s performance at a given threshold than other metrics Chicco & Jurman (2020, 2023). We encourage the community to adopt MCC for the epitope prediction task as it takes into account true and false positives, as well as true and false negatives, offering a comprehensive measure of the performance. It is considered superior to the AUC-ROC, which is evaluated over all thresholds. For consistency, we also included Precision and Recall from prior studies EpiPred Krawczyk et al. (2014) and PECAN Pittala & Bailey-Kellogg (2020), and we added Area Under the Receiver Operating Characteristic Curve (AUC-ROC) and F1 score, both are typical binary classification metrics. For methods that predict antibody-antigen complex structures, we determine the epitopes using the same distance criterion as aforementioned.
A.3 Steps to build antibody-antigen complex dataset
We sourced our initial dataset from AbDb (version dated September 26, 2022), containing 11,767 antibody files originally collected from the Protein Data Bank (PDB). We collected complexes numbered in Martin scheme Raghavan & Martin (2008) and used AbM CDR definition Martin et al. (1991) to identify CDR residues from the heavy and light chains of antibodies.
We extracted antibody-antigen complexes that met the following criteria: (1) both VH and VL domains are present in the antibody; (2) the antigen is a single-chain protein consisting of at least 50 amino acids; and (3) there are no unresolved CDR residues, yielding 4,081 files.
To deduplicate complexes, we used MMseqs2 Steinegger & Söding (2017) to cluster the complexes by heavy and light chains in antibodies and antigen sequences. We used the easy-linclust mode with the –cov-mode 0 option to cluster sequences; we used the default setting for coverage of aligned cutoff at 80%; we used different –min-seq-id cutoffs for antibodies and antigens because the antibody framework regions are more conserved than the CDR regions. We cluster heavy and light chains at –min-seq-id cutoff of 100% and 70%, respectively. We retained only one representative file for each unique set of identifiers, leading to a refined dataset of 1,725 files.
Two additional files were manually removed. File 6jmr_1P was removed because its CDR residues are masked with ‘UNK’ labels and the residue identities are unknown; file 7sgm_0P was removed because of a non-canonical residue ‘DV7’ in its CDR-L3 loop.
The final dataset consists of 1,723 antibody-antigen complexes.
A.4 Pipeline to build graph dataset from AbDb
Figure S1: Pipeline to convert an antibody-antigen complex structure into a graph representation.
  • Row 1: given an AbAg complex PDB ID, retrieve ‘AbAg complex’ from AbDb and ‘raw structure’ file from PDB as input, ‘AbDb complex antigen’ and ‘Original PDB file’ in the top lane.
  • Row 2: They are then parsed as hierarchical coordinates (Antigen Structure), and extract ATMSEQ and SEQRES sequences.
  • Row 3: these are then passed to a set of in-house modules for calculating solvent access surface area (ASA), distance matrix, and filtering out problematic residues, which generates an ATMSEQ2NODES mask. The sequence alignment module aligns ATMSEQ with SEQRES sequences to generate a mask mapping from SEQRES to ATMSEQ. The Sequence Embedding module passes SEQERS through the ESM module to generate embeddings. ESM requires input sequence length and therefore filters out sequences longer than 1021 amino acids.
  • Row 4: holds intermediate data that we apply masks to generate graph data in Row 5.
  • Row 6: Apply the masks to map SEQRES node embeddings to nodes in the graphs and calculate the edges between the graph nodes.
$U$, $V$ and $N$ denote the number of residues in the SEQRES sequence, ATMSEQ sequence and the graph, respectively. thr (at 50 residues) is the cutoff for antigen SEQRES length. We only include antigen sequences with lengths of at least 50 residues. SEQRES and ATMSEQ are two different sequence representations of a protein structure. SEQRES is the sequence of residues in the protein chain as defined in the header section of a PDB file, and it is the complete sequence of the protein chain. ATMSEQ is the sequence of residues in the protein chain as defined in the ATOM section of a PDB file. In other words, it is read from the structure, and any residues in a PDB file are not resolved due to experimental issues that will be missing in the ATMSEQ sequence. Since we are building graph representations using structures, we used ATMSEQ. However, the input to the language models require a complete sequence, therefore we used SEQRES to generate node embeddings, and mapped the node embeddings to the graph nodes. We performed two pairwise sequence alignments to map such embeddings to graph vertices for a protein chain with Clustal Omega Sievers et al. (2011). We first align the SEQRES sequence with the atom sequence (residues collected from the ATOM records in a PDB file) and assign residue embeddings to matched residues. Because we excluded buried residues from the graph, we aligned the sequence formed by the filtered graph vertices with the atom sequence to assign residue embeddings to vertices. ASA is the solvent-accessible surface area of a residue. If a residue has an ASA value of zero, it is considered buried and will be removed from the graph.
A.5 Dataset split
Table S3: Distribution of epitope to antigen surface nodes in each set.
Epi/Surf Training Validation/Test
0, 5% 320 (23% ) 40 (24%)
5% , 10% 483 (35% ) 60 (35%)
10%, 15% 305 (22% ) 38 (22%)
15%, 20% 193 (14% ) 24 (14%)
20%, 25% 53 (4% ) 6 (4% )
25%, 30% 19 (1% ) 2 (1% )
30%, 35% 8 (0.6%) 0 (- )
35%, 40% 2 (0.1%) 0 (- )
sum 1383 170
A.6 Implementation details
Exploratory Data Analysis We performed exploratory data analysis on the training dataset to understand the distribution of the number of residue-residue contacts in the antibody-antigen interface. We found that the number of contacts is approximately normally distributed with a mean of 43.42 and a standard deviation of 11.22 (Figure S2). We used this information to set the regularizer in the loss function to penalize the model for predicting too many or too few positive edges.
Figure S2: Blue line: distribution of the number of residue-residue contacts in antibody-antigen interface across the dataset with a mean and median of 43.27 and 43.00, respectively. Red line: fitted normal distribution with mean and standard deviation of 43.27 and 10.80, respectively.
Loss function Our loss function is a weighted sum of two parts: a binary cross-entropy loss for the bipartite graph linkage reconstruction and a regularizer for the number of positive edges in the reconstructed bipartite graph.
Loss $=\mathcal{L}_{r}+\lambda\left|\sum^{N}\hat{e}-c\right|$ (3)
$\mathcal{L}_{r}$ $=-\frac{1}{N}\sum_{i=1}^{N}\left(w_{\text{pos}}\cdot y_{e}\cdot\log(\hat{y_{e}})+w_{\text{neg}}\cdot(1-y_{e})\cdot\log(1-\hat{y_{e}})\right)$ (4)
The binary cross-entropy loss $\mathcal{L}_{r}$ is weighted by $w_{\text{pos}}$ and $w_{\text{neg}}$ for positive and negative edges, respectively. During hyperparameter tuning, we kept $w_{\text{neg}}$ fixed at $1.0$ and tuned $w_{\text{pos}}$. $N$ is the total number of edges in the bipartite graph, $y_{e}$ denotes the true label of edge $e$, and $\hat{y}_{e}$ denotes the predicted probability of edge $e$. The regularizer $\left|\sum^{N}\hat{e}-c\right|$ is the L1 norm of the difference between the sum of the predicted probabilities of all edges of the reconstructed bipartite graph and the mean positive edges in the training set, i.e., $c$ set to $43$. This aims to prevent an overly high false positive rate, given the fact that the number of positive edges is far less than positive edges. The regularizer weight $\lambda$ is tuned during hyperparameter tuning.
Hyperparameters We carried out hyperparameter search within a predefined space that included:
  • Weights for positive edges in bipartite graph reconstruction loss, sampled uniformly between 50 and 150.
  • Weights for the sum of bipartite graph positive links, where values were drawn from a log-uniform distribution spanning $71e-$ 7 to $41e-$ 4.
  • Edge cutoff (x), defining an epitope node as any antigen node with more than x edges, with x sampled following a normal distribution with a mean of 3 and a standard deviation of 1.
  • Number of graph convolutional layers in the encoder, we tested using 2 and 3 layers.
  • Decoder type was varied between two configurations:
    • A fully connected layer, equipped with a bias term and a dropout rate of 0.1.
    • An inner product decoder.
Computing resources All experiments were performed on a single Linux machine (Ubuntu 22.04) with one NVIDIA RTX3090 GPU card, and it took on average 3 hours for a single hyper-parameter sweeping experiment.
A.7 Antibody-antigen complex examples
To compare the epitopes of antibodies in AsEP, we first clustered these complexes by antigen sequences to group together antibodies targeting the same antigen via MMseqs2 Steinegger & Söding (2017). Specifically, we ran MMseqs2 on antigen SEQRES sequences and using the following setup:
  • easy-linclust mode
  • cov-mode set to 0 with the default coverage of 80%: this means a sequence is considered a cluster member if it aligns at least 80% of its length with the cluster representative;
  • min-seq-id set to 0.7: this means a sequence is considered a cluster member if it shares at least 70% sequence identity with the cluster representative.
We encourage the reader to refer to the MMseqs2 documentation https://github.com/soedinglab/mmseqs2/wiki for more details on the parameters used.
We then identify epitopes using a distance cut-off of 4.5 Å. An antigen residue is identified as epitope if any of its heavy atoms are located within 4.5 Å of any heavy atoms from the antibody.
To compare epitopes of antibodies sharing the same antigen cluster, we aligned the antigen SEQRES sequences using Clustal Omega Sievers et al. (2011) (download from: http://www.clustal.org/omega/) to obtain a Multiple Sequence Alignment (MSA). Epitopes are mapped to and denoted as the MSA column indices. The epitope similarity between a pair of epitopes is then calculated as the fraction of identical columns. Two epitopes are identified as identical if they share over 0.7 of identical columns.
Table S4: Antibody-Antigen Complex Examples
(a) Antigen Group Information
abdbid repr size epitope_group
7eam_1P 7sn2_0P 183 0
5kvf_0P 5kvd_0P 9 1
5kvg_0P 5kvd_0P 9 2
(b) CDR Sequences (Heavy Chain)
abdbid H1 H2 H3
7eam_1P GFNIKDTYIH RIDPGDGDTE FYDYVDYGMDY
5kvf_0P GYTFTSSWMH MIHPNSGSTN YYYDYDGMDY
5kvg_0P GYTFTSYGIS VIYPRSGNTY ENYGSVY
(c) CDR Sequences (Light Chain)
abdbid L1 L2 L3
7zf4_1P RASGNIHNYLA NAKTLAD QHFWSTPPWT
7zbu_0P KSSQSLLYSSNQKNYLA WASTRES QQYYTYPYT
7xxl_0P KASQNVGTAVA SASNRYT QQFSSYPYT
(d) Structure Titles
abdbid resolution title repr_title
7zf4_1P 1.4 immune complex of SARS-CoV-2 RBD and cross-neutralizing antibody 7D6 SARS-CoV-2 Omicron variant spike protein in complex with Fab XGv265
7zbu_0P 1.4 Zika specific antibody, ZV-64, bound to ZIKA envelope DIII Cryo-EM structure of zika virus complexed with Fab C10 at pH 8.0
7xxl_0P 1.4 Zika specific antibody, ZV-67, bound to ZIKA envelope DIII Cryo-EM structure of zika virus complexed with Fab C10 at pH 8.0
Here we provide three example antibody-antigen complexes from the same antigen group, meaning the antigen sequences from each member complex share sequence identity of the aligned region at least 70%. Due to space limitation, we have broken the rows into four parts: Antigen group information, CDR sequences, and structure titles.
  • abdbid: AbDb ID of the group member;
  • repr: AbDb ID of the antigen representative;
  • size: the number of complexes in the group;
  • epitope_group: A categorical identifier of the epitope group the antibody-antigen complex belongs to;
  • H1, H2, H3, L1, L2, L3: CDR sequences of the heavy and light chains;
  • resolution: Structure resolution;
  • title: Structure title of the member;
  • repr_title: Structure title of the antigen representative.
A.8 Multi-epitope Antigens
We include $641$ unique antigens and $973$ epitope groups in our dataset. Figure S3 shows two examples of multi-epitope antigens in our dataset, hen egg white lysozyme (Figure 3(a)) and spike protein (Figure 3(b)). Specifically, there are $52$ and $64$ distinct antibodies in our dataset that bind to hen egg white lysozyme and spike protein, respectively. For visual clarity, we only show five and sixteen antibodies in Figure 3(a) and Figure 3(b).
We can see that different antibodies bind to different locations on the same antigen. Details of all epitope groups are provided in the Supplementary Table SI-AsEP-entries.csv with annotation provided in Section A.7.
(a) Five different antibodies bound to hen egg white lysozyme. Complexes are superimposed on the antigen structure (magenta). AbDb IDs of the complexes and their color: 1g7i_0P (green), 2yss_0P (cyan), 1dzb_1P (yellow), 4tsb_0P (orange), 2iff_0P (wheat). Antigens are colored in magenta.
(b) Sixteen different antibodies bound to coronavirus spike protein. Complexes are superimposed on the antigen structure (magenta) and antibodies are in different colors. AbDb IDs of the complexes: 7k8s_0P, 7m7w_1P, 7d0b_0P, 7dzy_0P, 7ey5_1P, 7jv4_0P, 7k8v_1P, 7kn4_1P, 7lqw_0P, 7n8i_0P, 7q9i_0P, 7rq6_0P, 7s0e_0P, 7upl_1P, 7wk8_0P, 7wpd_0P.
Figure S3: Examples of different antibodies binding to the same antigen.
A.9 Metrics definition
$\mathrm{MCC}=\frac{(TP\times TN-FP\times FN)}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$
$\mathrm{Precision}=\frac{TP}{TP+FP}$
$\mathrm{Recall}=\frac{TP}{TP+FN}$
[…]
TP: True Positive
FP: False Positive
TN: True Negative
FN: False Negative
A.10 Fine-tuning ESMBind on AsEP
The performance reported in the main text for ESMBind is derived by fine-tuning ESM2 on general protein binding sites. We performed a further fine-tuning experiment, fine-tuning it on the presented AsEP dataset and evaluating it on the AsEP test set to enable a more direct comparison of ESMBind to WALLE. Fine-tuning of ESMBind on AsEP was done using the Low-Rank Adaptation method Hu et al. (2021).
Table S5: Performance Metrics
Metric Value
MCC 0.103923
Accuracy 0.504478
AUC-ROC 0.584497
Precision 0.128934
Recall 0.707731
F1 0.213829
Appendix B Appendix: Link prediction baseline
While the majority of existing studies focus on node-level prediction, i.e., predicting which residues are likely to be the epitope residues, we are interested in predicting the interactions between epitope and antigen residues. We argue that, on the one hand, this would provide a more comprehensive understanding of the interaction between epitopes and antigens, and on the other hand, it would be good in terms of model interpretability. Existing methods for predicting epitope residues are mostly based on sequence information, which is not directly interpretable in terms of the interaction between epitopes and antigens.
Our hyperparameter search was conducted within a predefined space as defined in Section A.6. We used the Bayesian optimization strategy implemented through Weights & Biases, targeting the maximization of the average bipartite graph link Matthew’s Correlation Coefficient (MCC).
The optimization process was managed using the early termination functionality provided by the Weights & Biases’ Hyperband method Falkner et al. (2018), with a range of minimum to maximum iterations set from 3 to 27.
The best set of hyperparameters is $2$ GCNConv layers, a batch size of $32$, a weight for positive edges of $54.7$, a weight for the sum of positive links at approximately $75.75e-$ 7, and an edge cutoff of $2.38$. The resulting MCC evaluated on the test set was $0.072$ (standard error: $0.009$) for the bipartite graph link prediction task.
Table S6: Evaluation of WALLE on the bipartite graph link prediction task
Metric Mean (Standard Error)
MCC 0.072 (0.009)
ROC-AUC 0.582 (0.011)
Precision 0.049 (0.008)
Recall 0.167 (0.023)
F1 0.053 (0.007)
Appendix C Appendix: Ablation Studies
To investigate the impact of different components on WALLE’s performance, we carried out ablation studies and described them in this section. For each model variant, we performed hyperparameter tuning and reported the evaluation performance using the model with the best performance on the validation set.
C.1 Ablation study: replace graph component with linear layers
To investigate whether the graph component within the WALLE framework is essential for its predictive performance, we conducted an ablation study in which the graph component was replaced with two linear layers. We refer to the model as ‘WALLE-L’. The first linear layer was followed by a ReLu activation function. Logits output by the second linear layer were used as input to the decoder. The rest of the model architecture remained the same.
It differs from the original WALLE model in that the input to the first linear layer is simply the concatenation of the embeddings of the antibody and antigen nodes, and the linear layers do not consider the graph structure, i.e., the spatial arrangement of either antibody or antigen residues. The model was trained using the same hyperparameters as the original WALLE model. The performance of WALLE-L was evaluated on the test set using the same metrics as the original WALLE model.
C.2 Ablation study: WALLE with simple node encoding
The presented WALLE model utilizes embeddings from large language models, including ESM2 Lin et al. (2022) or IgFoldRuffolo et al. (2023) for representing amino acid sequences, as these models are able to capture the sequential and structural information inherent in protein sequences, providing a rich, context-aware representation of amino acids. To test the effectiveness of such embeddings in this downstream task, we conducted an ablation study where we replaced the embeddings from language models with simple node encodings. Specifically, we evaluated the performance of WALLE when using ‘one-hot’ encoding and ‘BLOSUM62’ encoding for amino acids in both antibody and antigen sequences.
C.3 Ablation study: WALLE with ESM2 embeddings for both antibodies and antigens
We also investigated whether the choice of language models can impact the predictive performance of WALLE; we conducted an ablation study to evaluate the performance of WALLE when both antibodies and antigens are represented using embeddings from the ESM2 language model Lin et al. (2022) while the original model uses AntiBERTy Ruffolo et al. (2023) for antibodies as it is trained exclusively on antibody sequences. This also tests whether a language model trained on general protein sequences can be used for a downstream task like antibody-antigen interaction prediction.
One-hot encoding
One-hot encoding is a method where each residue is represented as a binary vector. Each position in the vector corresponds to a possible residue type, and the position corresponding to the residue present is marked with a 1, while all other positions are set to 0. This encoding scheme is straightforward and does not incorporate any information about the physical or chemical properties of the residues. This method tests the model’s capability to leverage structural and relational information from the graph component without any assumptions introduced by more complex encoding schemes.
BLOSUM62 encoding
BLOSUM62 Henikoff & Henikoff (1992) encoding involves using the BLOSUM62 matrix, which is a substitution matrix used for sequence alignment of proteins. In this encoding, each residue is represented by its corresponding row in the BLOSUM62 matrix. This method provides a more nuanced representation of residues, reflecting evolutionary relationships and substitution frequencies.
C.4 Hyperparameter tuning
We used the same hyperparameter search space defined in Section A.6 and performed a hyperparameter search as defined in Appendix B for each model variant in the ablation studies. We report the evaluation performance of the tuned model for each variant in Table S7.
Table S7: Performance of WALLE without graph component and simple node encodings on test set from dataset split by epitope to antigen surface ratio.
Algorithm Encoding MCC AUCROC Precision Recall F1
WALLE Both 0.2097 (0.0195) 0.6351 (0.0126) 0.2346 (0.0183) 0.4217 (0.0279) 0.2580 (0.0178)
WALLE-L Both 0.1593 (0.0155) 0.6124 (0.0109) 0.1750 (0.0109) 0.4696 (0.0243) 0.2371 (0.0137)
WALLE ESM2 0.1955 (0.0212) 0.6219 (0.0137) 0.2280 (0.0188) 0.4103 (0.0291) 0.2553 (0.0188)
WALLE-L ESM2 0.1445 (0.0138) 0.6100 (0.0097) 0.1598 (0.0102) 0.5355 (0.0216) 0.2266 (0.0125)
WALLE One-hot 0.0968 (0.0094) 0.5830 (0.0076) 0.1185 (0.0052) 0.8923 (0.0118) 0.2026 (0.0081)
WALLE BLOSUM 0.0848 (0.010) 0.5739 (0.0081) 0.1182 (0.0055) 0.8401 (0.0151) 0.1993 (0.0083)
The values in parentheses represent the standard error of the mean; ‘WALLE-L’ refers to WALLE with the graph component replaced by two linear layers. ‘ESM2’ refers to the embeddings from the ESM2 language model esm2_t12_35M_UR50D. ‘One-Hot’ refers to one-hot encoding of amino acids. ‘BLOSUM62’ refers to the BLOSUM62 encoding of amino acids. ‘Both’ refers to embedding antibodies and antigens using the esm2_t12_35M_UR50D ESM2 model and AntiBERTy (via IgFold) language model, respectively. The best performing model is highlighted in bold.
We observed that WALLE’s performance with simple node encodings (‘one-hot’ and ‘BLOSUM62’) is considerably lower than when using advanced embeddings from language models. This indicates that the embeddings derived from language models capture more nuanced information about the amino acids, enabling the model to better predict epitope-antigen interactions.
The degenerated performance of WALLE with simple encodings can be attributed to the lack of contextual information and structural features in these representations. The high recall but low precision values suggest that the model is unable to distinguish between true and false interactions, leading to a high number of false positives. This highlights the importance of using meaningful embeddings that capture the rich structural and sequential information present in protein sequences.
When comparing WALLE with WALLE-L (without the graph components), we observe that the model’s performance drops considerably when the graph component is replaced with fully connected linear layers. This indicates that the topological information captured by the graph component also contributes to the model’s predictive performance.
We also observed that WALLE with ESM2 embeddings for both antibodies and antigens achieved similar performance to WALLE with AntiBERTy and ESM2 embeddings for antibodies and antigens, respectively. This suggests that the ESM2 embeddings somehow provide effective information for both antibodies and antigens without training exclusively on antibody sequences.
Appendix D Impact Statement
This work extends to the optimization of antibody drug development, offering a more efficient and accurate method for predicting where antibodies bind on antigens, a crucial challenge of developing therapeutic antibodies. The potential impact of this advancement is underscored by the recent approval of 136 antibody-based drugs in the United States or European Union, with 17 novel antibody therapeutics approved since January 2023 and 18 more currently under review (Antibody Society, Antibody therapeutics approved or in regulatory review in the EU or US, https://www.antibodysociety.org/resources/approved-antibodies/, 24 January 2024). These figures show the transformative impact that antibody design innovation can have on transforming the landscape of therapeutic development, offering new avenues for targeted, effective treatments that can be developed in a more time- and cost-effective manner. Such advancements hold immense potential in accelerating the creation of targeted therapies and implicitly support the broader goals of personalized medicine. Fast-tracking antibody design through in silico epitope prediction can enable quicker responses to emerging health threats like pandemics or rapidly mutating pathogens. However, in silico predictions should not be blindly trusted and should merely guide and streamline research and diligent testing, not replace it.
This paper presents work whose goal is to highlight antibody-specific epitope prediction as a major challenge that has not been widely studied to date, which is presented as a bipartite graph connectivity prediction task. This research stands out primarily for its development of a novel benchmarking dataset for antibody-specific epitope prediction - a resource previously unavailable in the scientific community with the potential to set a new standard in the field.
Existing methods are compared using uniform metrics and a vast room for improvement can be demonstrated across the board. A novel model is presented, which surpasses its counterparts by an order of magnitude, outperforming them by a factor of ten, validating its graph-based approach utilizing antibody and antigen structural information as well as leveraging protein language models. We provide a foundational framework that invites other researchers to further build upon and refine this baseline model, which we have demonstrated to be a highly effective approach for this task.
Open availability of the dataset and model facilitates further research and exploration in this domain, expediting the development of more advanced models. Highlighting types of antibody-antigen interactions that are disregarded in existing datasets and methods encourages the examination of shortcoming of current models. The focus on conventional antibodies in this work lays the groundwork for future exploration into epitope prediction for novel antibodies, such as nanobodies, expanding upon potential applications.
Supplementary Materials
Dataset Documentation and Intended Uses
We provide a data card for this dataset, DataCard-AsEP.md, which can be downloaded using this link: https://drive.google.com/file/d/1fc5kFcmUdKhyt3WmS30oLLPgnkyEeUjJ/view?usp=drive_link
This dataset provides a unified benchmark for researchers to develop new machine-learning-based methods for the epitope prediction task.
Access to the Dataset
There are two alternative sources where users can download the dataset:
  • The dataset can be downloaded using the Python interface provided by our GitHub Repository AsEP-dataset. Detailed instructions on how to download the dataset are provided in the README file. Briefly, after installing the provided Python module, asep, the dataset can be downloaded by running the following command in the terminal:
    download-asep /path/to/directory AsEP
    # For example, to download the dataset to the current directory, run
    # download-asep . AsEP
  • The dataset and benchmark are provided through Zenodo at https://doi.org/10.5281/zenodo.11495514.
  • Code and Dataset interface is provided in our GitHub Repository AsEP-dataset at https://github.com/biochunan/AsEP-dataset.
Author Statement
The authors affirm that they bear all responsibility in case of violation of rights, etc., and confirm the data license. The dataset is licensed under the CC BY 4.0 License (https://creativecommons.org/licenses/by/4.0/), which is provided through the Zenodo repository. The code is licensed under the MIT License (https://opensource.org/licenses/MIT).
Hosting, Licensing, and Maintenance Plan
The dataset is hosted on Zenodo, which provides a DOI (10.5281/zenodo.11495514) for the dataset. It also comes with a Python interface provided in our GitHub Repository, AsEP-dataset at https://github.com/biochunan/AsEP-dataset, where users can submit issues and ask questions. Future releases and updates will be made available through the same channels. As discussed in the main text, the future plan includes expanding the dataset to include novel types of antibodies, such as single-domain antibodies, and providing more sophisticated features for graph representations. The dataset will be maintained by the authors and will be available for a long time.
Links to Access the Dataset and Its Metadata
The dataset, benchmark, and metadata are provided through Zenodo.
The Dataset
The dataset is constructed using pytorch-geometric Dataset module. The dataset can be loaded using the following code:
from asep.data.asepv1_dataset import AsEPv1Evaluator
evaluator = AsEPv1Evaluator()
# example
torch.manual_seed(0)
y_pred = torch.rand(1000)
y_true = torch.randint(0, 2, (1000,))
input_dict = {’y_pred’: y_pred, ’y_true’: y_true}
result_dict = evaluator.eval(input_dict)
print(result_dict) # got {’auc-prc’: tensor(0.5565)}
We also provide detailed documentation of the dataset content on Zenodo and include a description below:
  • asepv1-AbDb-IDs.txt: A text file containing the AbDb identifiers of the 1723 antibody-antigen pairs in the dataset.
  • asepv1_interim_graphs.tar.gz: Contains 1723 .pt files, where each file is a dictionary with structured data:
    abdbid
    A string representing the antibody AbDb identifier.
    seqres
    A dictionary containing:
    ab
    An OrderedDict mapping string chain labels H and L to their corresponding sequence strings, representing heavy and light chains respectively.
    ag
    A dictionary mapping string chain labels to their corresponding sequence strings.
    mapping
    Includes:
    ab
    Contains:
    seqres2cdr
    A binary numpy array indicating the CDR positions in the antibody sequence.
    ag
    Contains:
    seqres2surf
    A binary numpy array indicating the surface residues in the antigen sequence.
    seqres2epitope
    A binary numpy array indicating the epitope residues in the antigen sequence.
    embedding
    Comprises:
    ab
    Includes embeddings computed using the AntiBERTy model and ESM2 model for the antibody sequences.
    ag
    Includes embeddings for the antigen sequences computed using the ESM2 model.
    edges
    Describes the interactions:
    ab
    A sparse coo tensor representing the binary edges between the CDR residues.
    ag
    A sparse coo tensor representing the binary edges between the surface residues.
    stats
    Metadata about each antibody-antigen pair, including counts of CDR, surface, and epitope residues, and the epitope-to-surface ratio.
  • structures.tar.gz: Contains 1723 pdb structures, each named using the AbDb identifier.
  • split_dict.pt: Contains the train/val/test splits of the dataset, with splits based on the epitope ratio and epitope group of the antigen.
Long-term Preservation
The current version of the dataset and benchmark are provided through Zenodo, which provides long-term storage. Future versions will be made available through the same channel and users are encouraged to submit queries and issues through the issues channel on the GitHub repository.
Explicit License
The dataset is licensed under the CC BY 4.0 license (https://creativecommons.org/licenses/by/4.0/), which is provided through the Zenodo repository. The code is licensed under the MIT License https://opensource.org/licenses/MIT.
Benchmarks
Detailed benchmark experiments and results are provided on Zenodo (https://doi.org/10.5281/zenodo.11495514), and the file benchmark.zip contains the instructions on how to reproduce the results. To run ESMFold, EpiPred, and MaSIF-Site, we provided docker images on the Zenodo repository or instructions on how to obtain them from DockerHub. The benchmark results are reproducible by following the instructions provided in the zip file. For the method WALLE and its variants used in the ablation studies, their configuration YAML files and the trained models are also provided on Zenodo.